OpenStudy (anonymous):

(-9x^2) ÷ (x^2 - 4x^3)

OpenStudy (anonymous):

= x^2(-9)/[x^2(1-4x)] = -9/(1-4x)

OpenStudy (anonymous):

or.. 9/(4x-1)